### v1

# LLMs in Biomedicine: A study on clinical Named Entity Recognition

Abstract
Large Language Models (LLMs) demonstrate remarkable versatility in various NLP tasks but encounter distinct challenges in biomedicine due to medical language complexities and data scarcity. This paper investigates the application of LLMs in the medical domain by exploring strategies to enhance their performance for the Named-Entity Recognition (NER) task. Specifically, our study reveals the importance of meticulously designed prompts in biomedicine. Strategic selection of in-context examples yields a notable improvement, showcasing $\sim 15-20\%$ increase in F1 score across all benchmark datasets for few-shot clinical NER. Additionally, our findings suggest that integrating external resources through prompting strategies can bridge the gap between general-purpose LLM proficiency and the specialized demands of medical NER. Leveraging a medical knowledge base, our proposed method inspired by Retrieval-Augmented Generation (RAG) can boost the F1 score of LLMs for zero-shot clinical NER. We will release the code upon publication.
LLMs in Biomedicine: A study on clinical Named Entity Recognition
Masoud Monajatipoor UCLA monajati@ucla.edu                        Jiaxin Yang UCLA yangjiaxin0821@ucla.edu                        Joel Stremmel Optum joel_stremmel@optum.com
Melika Emami Optum melika.emami@optum.com                        Ehsan Mohaghegh Optum ehsan.mohaghegh@optum.com                        Mozhdeh Rouhsedaghat USC rouhseda@usc.edu                        Kai-Wei Chang UCLA kwchang@cs.ucla.edu
1 Introduction
LLMs such as OpenAI’s GPT-4 have demonstrated exceptional capabilities across diverse tasks and domains, exhibiting remarkable performance Espejel et al. (2023); Dai et al. (2023); Dong et al. (2022). These models could have a revolutionary impact on healthcare, however, their integration into medical research and practice lags behind Zhou et al. (2023); Vaishya et al. (2023); Nori et al. (2023a). Acknowledging this gap, it is crucial to examine the unique challenges presented by the medical field that contribute to this discrepancy. Specifically, LLMs encounter challenges in medical Information Extraction (IE) Gutierrez et al. (2022); Moradi et al. (2021) due to the scarcity of high-quality clinical data in their pretraining, compounded by the requirement for a nuanced comprehension of the text for this task Gu et al. (2023). Medical entities can have multiple synonyms and abbreviations which makes them difficult for models to recognize Grossman Liu et al. (2021). Besides, context sensitivity takes on even greater significance in biomedicine compared to the general domain. The specificity of entity types and the complexity of their interrelations necessitate a level of background knowledge that standard prompts may fail to provide. LLMs are primarily exposed to vast amounts of generic text data limiting their effectiveness in managing the intricate nuances of medical language Kumari et al. (2023); Karabacak and Margetis (2023).
In this paper, we concentrate on NER, a foundational task for various applications in biomedicine. In our initial analysis, we broaden the scope of TANL Paolini et al. (2021) and DICE Ma et al. (2022), two text-to-text formats initially proposed for model training, adapting their use to prompt design specifically for the biomedical NER. Our findings reveal that the relative effectiveness of the resulting prompt pattern varies based on specific dataset characteristics. Subsequently, we investigate the importance of example selection in In-Context Learning (ICL) and demonstrate the value of nearest neighbor example selection using example encoders pretrained on biomedical text for the biomedical NER. A key question that arises in the deployment of LLMs concerns the comparative advantage of closed-source LLMs versus fine-tuned open-source ones. In our third study, we shed light on this question by presenting an assessment of performance and cost across various experiments. Furthermore, we explore the integration of external clinical knowledge to further refine LLM capabilities Gao et al. (2023c); Zakka et al. (2024). Leveraging the insights gained from these techniques, we present a novel data augmentation method incorporating a medical knowledge base, e.g., UMLS Bodenreider (2004), which substantially improves zero-shot clinical NER.
2 Background and Preliminaries
Prompt engineering
Prompt tuning White et al. (2023); Lester et al. (2021); Ding et al. (2021) as its own research field shows that skillfully crafted prompts can significantly enhance LLM understanding for complex tasks Lu et al. (2021); Kaddour et al. (2023); Webson and Pavlick (2021). Researchers have explored different formats to examine how LLMs handle IE tasks Wang et al. (2023c); Gutierrez et al. (2022); Wang et al. (2023b). Another type of prompting is ICL Brown et al. (2020), where LLMs use a limited set of "input-output" pairs within the prompt along with a query input as demonstrations of what the task output should be. In this realm, Liu et al. (2021); Min et al. (2022); Gao et al. (2023a) demonstrated that choosing targeted in-context examples over random sampling leads to more accurate model responses.
Named Entity Recognition
GPT-NER Wang et al. (2023b) was one of the first methods to incorporate a unique symbol to transform the sequence tagging task into text generation via ICL with GPT-3 Brown et al. (2020), achieving performance levels on par with fully supervised baselines. Following this work, Gutierrez et al. (2022); Moradi et al. (2021) have shown that LLMs are not skilled few-shot learners in the biomedical domain. However, recent advancements, such as GPT-4, have increased LLM performance on various tasks Tian et al. (2024); Hu et al. (2024a); Nori et al. (2023a) including in the biomedical domain Hu et al. (2024b). In the direction of knowledge distillation from LLMs Wang et al. (2023c); Gu et al. (2023), UniNER Zhou et al. (2023) presented a targeted distillation technique coupled with instruction tuning to develop an efficient open-domain NER model. Our research draws from these works and explores the capabilities of LLMs for clinical NER, employing prompt design, strategic ICL example selection, and data augmentation via an external knowledge base to enhance performance.
Problem definition
Assume the data sample is represented as $(X,Y)$ and the goal is to develop a model, denoted as $f:(X\times T)\rightarrow Y$, where X signifies the input set, T represents a predetermined collection of entity types, and Y denotes the set of entities types within the provided input. The task is to predict the entity type of each input word among the set T. We followed the standard practice of using the F1 score for evaluation purposes.
Datasets
We use three biomedical NER datasets with different entity types: I2B2 Uzuner et al. (2011) which includes test, treatment, and problem entities, NCBI-disease Doğan et al. (2014) consisting of the disease entity, and BC2GM Smith et al. (2008) containing the gene entity.
3 Influence of Input-Output Format
Recent studies demonstrated the importance of prompt engineering for various tasks Wang et al. (2023a); Gao et al. (2023b); Nori et al. (2023b). We studied the influence of input-output format by adapting TANL Paolini et al. (2021) and DICE Ma et al. (2022) formats for the biomedical NER. In TANL, the task is framed as a translation task which involves augmenting the text by tagging entity types for each word directly within the text. The method is exemplified in Fig 7, showcasing how the text incorporates entity types.
[…]
Figure 1: TANL input/output format for NER task.
Then, the generated output is decoded into the BIO format for the assessment. In the refined DICE format, the input-output format involves adding a description for each entity type in a template following DEGREE Hsu et al. (2021). Given an input text and corresponding labels, the desired output should be the input followed by the phrase "entity type is <entity_type>. <entity_description>. entity is <entity>" for each class label, e.g., test, treatment, and problem in the I2B2 dataset. Then, we expect the model to output the same template filling out the <entity> with the corresponding entities in the given text as demonstrated in Fig 2. For the entity types with no matched entities in the sentence, the output returns <entity> token in the output. Examples for the NCBI-disease and BC2GM datasets are presented in Appendix A.
[…]
Figure 2: DICE input/output format for NER task.
Our experiments in Table 1 reveal that neither format consistently outperforms the other, rather the effectiveness of each format varies depending on the complexity of the dataset and model size. To maintain consistency in the rest of our experiments, we opted for the TANL format, in which the input-output relationship exhibits a more straightforward pattern.
input-output format I2B2 NCBI-disease BC2GM
GPT-3.5-turbo DICE 50.0 62.0 55.6
TANL 59.7 51.3 50.8
GPT-4 DICE 70.1 77.8 67.9
TANL 73.5 70.0 69.6
Table 1: Comparison of TANL and DICE using GPT-3.5-turbo/GPT-4 on three benchmark datasets. The superiority of any single format varies with the complexity of the dataset and model size.
4 In-Context Examples Selection: A Key to Improving ICL Outcomes
In-context examples can be randomly chosen from the training set, however, researchers have demonstrated that the performance of ICL depends on the order and similarity of ICL examples to the test samples Liu et al. (2021); Min et al. (2022); Gao et al. (2023a); Zhang et al. (2023). Liu et al. (2021) presented Knn-Augmented in-conText Example selection (KATE). KATE identifies in-context examples selectively using nearest neighbor search on example embeddings, leading to better performance than random example selection. We tested KATE on TANL formatted examples with 16-shot ICL using four different LM encoders (w/o fine-tuning) to produce example embeddings. We used MPNET Song et al. (2020) for its popularity and performance on sentence embedding benchmarks111 https://www.sbert.net/docs/pretrained_models.html, SimCSE Gao et al. (2021) for its documented performance as an alternative to standard sentence transformers, and BioClinicalBERT Alsentzer et al. (2019) and BioClinicalRoBERTa Gururangan et al. (2020) for their dominance on clinical data tasks Lehman et al. (2023).
Our results summarized in Table 2 show that strategic in-context example selection via KATE outperforms random selection. BioClinicalRoBERTa achieved the best results among all example encoders tested. The strong performance of BioClinicalBERT and BioClinicalRoBERTa underscores the importance of using LM encoders pretrained on the biomedical text when applying KATE for the biomedical NER.
KATE method vs random sampling I2B2 NCBI-disease BC2GM
GPT-3.5-turbo Random sampling 59.7 51.3 50.8
[…]
BioClinicalBERT 78.9 78.8 72.0
[…]
SimCSE 76.1 76.1 68.8
Hu et al. (2024b) 49.3 - -
GPT4 Random sampling 73.5 70.0 69.6
[…]
BioClinicalBERT 88.1 88.0 79.4
[…]
SimCSE 86.6 86.5 77.9
Hu et al. (2024b) 59.3 - -
Table 2: Random example selection versus KATE with medical/non-medical encoders. KATE significantly outperforms random sampling in all settings, and LMs pretrained on the biomedical text outperform strong, general domain encoders.
[…]
Figure 3: An overview of Dictionary-Infused RAG
5 In-Context Learning or Fine-Tuning?
Within the scope of LLMs for biomedical applications, an essential question is whether to prompt a closed-source LLM via ICL or fine-tune an open-source one. On the other hand, comparing two different LLMs employing divergent strategies is not straightforward. To provide some insight into this dilemma, we examined two key factors, performance and cost, for clinical NER, and present a detailed analysis under various experiment settings. This comparison offers valuable perspective into the right strategy given the task and dataset attributes. For fine-tuning, we used LoRA Hu et al. (2021), and its details can be found in Appendix A. The cost of fine-tuning comes from training an LLM on a large labeled dataset while the cost of ICL mainly comes from calling the API for each input query. For 16-shot ICL experiments, we calculated the cost based on the number of processed and generated tokens considering the average text size based on the current LLM API pricing.222 https://openai.com/pricing The estimated cost for the entire test set of each benchmark dataset considering the input text, prompt, and generated text size using the TANL format is summarized in Table 3. Referring to the OpenAI API for fine-tuning pricing, we also estimated the cost for fine-tuning LLama2-7B which is summarized in Table 3. Interestingly, for the I2B2 dataset, GPT-3.5-turbo with a much cheaper cost outperforms fine-tuning Llama2-7B.
Language model I2B2 NCBI-disease BC2GM
Performance GPT-3.5-turbo w/ KATE 78.9 78.8 72.5
GPT4 w/ KATE 88.4 88.3 80.7
Llama2-7B 76.2 91.3 75.1
Cost GPT3.5-turbo w/ KATE \$0.35 \$0.11 \$1.34
GPT4 w/ KATE \$10.42 \$3.12 \$40.13
Llama2-7B \$5.23 \$5.24 \$5.69
Table 3: Analysis of ICL vs fine-tuning LLMs: assessing performance and cost implications. Fine-tuning LLama2 exhibits superior outcomes on NCBI-disease, whereas GPT-4, enhanced by KATE using a biomedical encoder, achieves more favorable results on both the I2B2 and BC2GM datasets.
6 Dictionary-Infused RAG
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a technique to enhance the capabilities of LLMs by integrating external information or knowledge into the generation process. This method involves retrieving relevant documents from a large corpus and providing this external knowledge in the input context to improve the quality and relevance of the generated text. Inspired by RAG, we developed a new method, DiRAG, to utilize UMLS as an external resource to augment the input data for the clinical NER task. The process with detailed prompts is visualized in Fig 3. Unlike traditional RAG techniques that rely on embedding similarities to retrieve relevant documents, our approach initially employs GPT-3.5-turbo to tackle a more straightforward task: identifying all words within an input text that could potentially qualify as named entities. Then, we look up each selected word in an external knowledge base, e.g., UMLS to augment the input data with useful information such as definition. Finally, we call GPT-4 with augmented input text. A similar prompting format to UniNER is used to do a fair comparison. We tested this approach on zero-shot NER and compared it with baselines in Table 4. Our proposed strategy with GPT-4 outperforms the baseline on I2B2 and NCBI-disease datasets with an F1 score improvement of 5.6% and 0.5% respectively. Additionally, the approach enhanced the performance of GPT-3.5-turbo on both the I2B2 and NCBI-disease datasets by 10.4% and 6.9% respectively. It proved ineffective for the BC2GM dataset due to the nature of the UMLS knowledge base which is predominantly tailored to medical terminology rather than biogenetics. We expect our approach to outperform GPT-4 on BC2GM if a more relevant knowledge base is employed.
I2B2 NCBI-disease BC2GM
UniversalNER Zhou et al. (2023) 40.4 61.4 47.2
Rohanian et al. (2023) w/ GPT-3.5 - 33.4 32.0
Hu et al. (2024b) w/ GPT-3.5-turbo 43.6 - -
Hu et al. (2024b) w/ GPT-4 52.6 - -
[…]
GPT-3.5-turbo w/ DiRAG 43.55 45.0 28.8
[…]
GPT-4 w/ DiRAG 60.04 62.17 53.83
Table 4: Zero-shot clinical NER using GPT-3.5-turbo and GPT-4 w/ and w/o DiRAG in comparison to SOTA. Our method improved zero-shot NER significantly for I2B2 and NCBI-disease datasets.
7 Conclusion
In this paper, we explored LLMs by customizing various prompting techniques specifically for the biomedical NER. Through a detailed comparative analysis, we have highlighted the vital role of ICL and the selection of contextually pertinent examples, demonstrating the benefits to clinical NER tasks. Moreover, our investigation into incorporating external clinical knowledge resulted in a novel data augmentation approach, considerably advancing the capabilities of zero-shot clinical NER with LLMs.
Limitations
While we have shown the potential of enhancing LLM performance for clinical NER, the experiments in this paper are limited in two aspects mainly due to computational constraints. (1) TANL uses a straightforward text-to-text format while DICE uses additional descriptions. Future work could attempt to simplify DICE or combine it with TANL. Ablation studies on components of each format could help researchers design new prompt formatting strategies. (2) Our RAG-based method exclusively utilizes UMLS as the knowledge base which is limited in its vocabulary. For medical terms not covered by UMLS, we did not augment the input text, though other knowledge bases such as Wikipedia could potentially serve as an alternative.
8 Aknowledgment
This work was supported in part by Optum. We would like to thank the reviewers for their feedback to improve this research work.
Appendix A Appendix
Fine-tuning setting for Llama
We fine-tuned Llama2-7B for three epochs and maintained a batch size of 16, learning rate of 2e-4, and cap the maximum sequence length at 512, truncating any sequences that exceeded this limit. The LoRA dropout rate is adjusted to 0.1, and the LoRA $\alpha$ and rank parameters are also set at 16 and 32 respectively. The training was done on 4 NVIDIA Tesla V100 GPUs for an approximate of 2 days.
TANL and DICE more examples
In this section, we visualize two examples of each format for NCBI-disease and BC2GM datasets for demonstration.
[…]
Figure 4: TANL input-output format example for NCBI-disease dataset
[…]
Figure 5: DICE input-output format example for NCBI-disease dataset
[…]
Figure 6: TANL input-output format example for BC2GM dataset
[…]
Figure 7: DICE input-output format example for BC2GM dataset

### v2

@@
-LLMs in Biomedicine: A study on clinical Named Entity Recognition
+LLMs in Biomedical: A Study on Named Entity Recognition
 Abstract
-Large Language Models (LLMs) demonstrate remarkable versatility in various NLP tasks but encounter distinct challenges in biomedicine due to medical language complexities and data scarcity.
-This paper investigates the application of LLMs in the medical domain by exploring strategies to enhance their performance for the Named-Entity Recognition (NER) task.
-Specifically, our study reveals the importance of meticulously designed prompts in biomedicine.
-Strategic selection of in-context examples yields a notable improvement, showcasing $\sim 15-20\%$ increase in F1 score across all benchmark datasets for few-shot clinical NER.
-Additionally, our findings suggest that integrating external resources through prompting strategies can bridge the gap between general-purpose LLM proficiency and the specialized demands of medical NER.
-Leveraging a medical knowledge base, our proposed method inspired by Retrieval-Augmented Generation (RAG) can boost the F1 score of LLMs for zero-shot clinical NER.
-We will release the code upon publication.
-LLMs in Biomedicine: A study on clinical Named Entity Recognition
+Large Language Models (LLMs) demonstrate remarkable versatility in various NLP tasks but encounter distinct challenges in biomedical due to the complexities of language and data scarcity.
+This paper investigates LLMs application in the biomedical domain by exploring strategies to enhance their performance for the NER task.
+Our study reveals the importance of meticulously designed prompts in the biomedical.
+Strategic selection of in-context examples yields a marked improvement, offering $\sim 15-20\%$ increase in F1 score across all benchmark datasets for biomedical few-shot NER.
+Additionally, our results indicate that integrating external biomedical knowledge via prompting strategies can enhance the proficiency of general-purpose LLMs to meet the specialized needs of biomedical NER.
+Leveraging a medical knowledge base, our proposed method, DiRAG, inspired by Retrieval-Augmented Generation (RAG), can boost the zero-shot F1 score of LLMs for biomedical NER.
+Code is released at https://github.com/masoud-monajati/LLM_Bio_NER
+LLMs in Biomedical: A Study on Named Entity Recognition
 Masoud Monajatipoor UCLA monajati@ucla.edu                        Jiaxin Yang UCLA yangjiaxin0821@ucla.edu                        Joel Stremmel Optum joel_stremmel@optum.com
-Melika Emami Optum melika.emami@optum.com                        Ehsan Mohaghegh Optum ehsan.mohaghegh@optum.com                        Mozhdeh Rouhsedaghat USC rouhseda@usc.edu                        Kai-Wei Chang UCLA kwchang@cs.ucla.edu
+Melika Emami Optum melika.emami@optum.com                        Fazlolah Mohaghegh Optum ehsan.mohaghegh@optum.com                        Mozhdeh Rouhsedaghat USC rouhseda@usc.edu                        Kai-Wei Chang UCLA kwchang@cs.ucla.edu
 1 Introduction
-LLMs such as OpenAI’s GPT-4 have demonstrated exceptional capabilities across diverse tasks and domains, exhibiting remarkable performance Espejel et al.
+LLMs such as GPT4 have demonstrated exceptional capabilities across diverse tasks and domains Espejel et al.
 (2023); Dai et al.
@@
 (2022).
-These models could have a revolutionary impact on healthcare, however, their integration into medical research and practice lags behind Zhou et al.
+These models could have a revolutionary impact on healthcare; however, their integration into medical research and practice has been slow Zhou et al.
 (2023); Vaishya et al.
 (2023); Nori et al.
-(2023a).
-Acknowledging this gap, it is crucial to examine the unique challenges presented by the medical field that contribute to this discrepancy.
-Specifically, LLMs encounter challenges in medical Information Extraction (IE) Gutierrez et al.
+(2023a) and it is crucial to examine the unique challenges presented by the biomedical field that contribute to this discrepancy.
+Specifically, LLMs encounter challenges in medical Information Extraction Gutierrez et al.
 (2022); Moradi et al.
-(2021) due to the scarcity of high-quality clinical data in their pretraining, compounded by the requirement for a nuanced comprehension of the text for this task Gu et al.
+(2021) due to the scarcity of high-quality biomedical data in their pretraining, and the need for a nuanced comprehension of the text for this task Gu et al.
 (2023).
-Medical entities can have multiple synonyms and abbreviations which makes them difficult for models to recognize Grossman Liu et al.
+Medical entities can have multiple synonyms and abbreviations, complicating their recognition by models Grossman Liu et al.
 (2021).
-Besides, context sensitivity takes on even greater significance in biomedicine compared to the general domain.
+Furthermore, context sensitivity is even more critical in the biomedical compared to the general domain.
 The specificity of entity types and the complexity of their interrelations necessitate a level of background knowledge that standard prompts may fail to provide.
@@
 (2023); Karabacak and Margetis (2023).
-In this paper, we concentrate on NER, a foundational task for various applications in biomedicine.
+In this paper, we concentrate on NER, a foundational task for various applications such as recruiting patients for clinical trials, searching biomedical literature, or building models that predict the progression of disease based on free-text notes.
 In our initial analysis, we broaden the scope of TANL Paolini et al.
 (2021) and DICE Ma et al.
-(2022), two text-to-text formats initially proposed for model training, adapting their use to prompt design specifically for the biomedical NER.
+(2022), two text-to-text formats initially proposed for model training, adapting their use to prompt design specifically for biomedical NER.
 Our findings reveal that the relative effectiveness of the resulting prompt pattern varies based on specific dataset characteristics.
-Subsequently, we investigate the importance of example selection in In-Context Learning (ICL) and demonstrate the value of nearest neighbor example selection using example encoders pretrained on biomedical text for the biomedical NER.
-A key question that arises in the deployment of LLMs concerns the comparative advantage of closed-source LLMs versus fine-tuned open-source ones.
+Subsequently, we investigate the importance of example selection via In-Context Learning (ICL) and demonstrate the value of nearest neighbor example selection using pre-trained biomedical text encoders when performing biomedical NER.
+A key question that arises in the deployment of LLMs concerns the comparative advantage of closed-source LLMs versus open-source ones.
 In our third study, we shed light on this question by presenting an assessment of performance and cost across various experiments.
-Furthermore, we explore the integration of external clinical knowledge to further refine LLM capabilities Gao et al.
+Furthermore, we explore the integration of external medical knowledge to refine LLM capabilities Gao et al.
 (2023c); Zakka et al.
 (2024).
-Leveraging the insights gained from these techniques, we present a novel data augmentation method incorporating a medical knowledge base, e.g., UMLS Bodenreider (2004), which substantially improves zero-shot clinical NER.
+Leveraging the insights gained from these techniques, we present a novel data augmentation method incorporating a medical knowledge base, e.g., UMLS Bodenreider (2004), which substantially improves zero-shot biomedical NER.
 2 Background and Preliminaries
@@
 (2023); Webson and Pavlick (2021).
-Researchers have explored different formats to examine how LLMs handle IE tasks Wang et al.
+Researchers have explored different prompt formats for IE tasks with LLMs Wang et al.
 (2023c); Gutierrez et al.
 (2022); Wang et al.
-(2023b).
-Another type of prompting is ICL Brown et al.
+(2023b) including more work around knowledge insertion for prompt augmentation Seo et al.
+(2024); Chen et al.
+(2023) Another type of prompting is ICL Brown et al.
 (2020), where LLMs use a limited set of "input-output" pairs within the prompt along with a query input as demonstrations of what the task output should be.
@@
 (2023b) was one of the first methods to incorporate a unique symbol to transform the sequence tagging task into text generation via ICL with GPT-3 Brown et al.
-(2020), achieving performance levels on par with fully supervised baselines.
+(2020), achieving performance on par with fully supervised baselines.
 Following this work, Gutierrez et al.
 (2022); Moradi et al.
-(2021) have shown that LLMs are not skilled few-shot learners in the biomedical domain.
-However, recent advancements, such as GPT-4, have increased LLM performance on various tasks Tian et al.
+(2021) showed that LLMs are not skilled few-shot learners in the biomedical domain.
+However, recent advancements, such as GPT-4, have increased LLM performance on many tasks Tian et al.
 (2024); Hu et al.
@@
 (2023c); Gu et al.
-(2023), UniNER Zhou et al.
-(2023) presented a targeted distillation technique coupled with instruction tuning to develop an efficient open-domain NER model.
-Our research draws from these works and explores the capabilities of LLMs for clinical NER, employing prompt design, strategic ICL example selection, and data augmentation via an external knowledge base to enhance performance.
+(2023), Zhou et al.
+(2023) presented UniNER, a targeted distillation technique coupled with instruction tuning to develop an efficient open-domain NER model.
+Our research draws from these works and explores the capabilities of LLMs for biomedical NER, employing prompt design, strategic ICL example selection, and data augmentation via an external knowledge base to enhance performance.
 Problem definition
-Assume the data sample is represented as $(X,Y)$ and the goal is to develop a model, denoted as $f:(X\times T)\rightarrow Y$, where X signifies the input set, T represents a predetermined collection of entity types, and Y denotes the set of entities types within the provided input.
+Assume data samples are represented as $(X,Y)$ and the goal is to develop a model, denoted as $f:(X\times T)\rightarrow Y$, where X signifies the input set, T represents a predetermined set of entity types, and Y denotes the set of entity types.
 The task is to predict the entity type of each input word among the set T.
-We followed the standard practice of using the F1 score for evaluation purposes.
+We followed the standard practice of using the F1 score for evaluation purposes in both mention/token-level analyses.
 Datasets
-We use three biomedical NER datasets with different entity types: I2B2 Uzuner et al.
+We used three biomedical NER datasets with different entity types: I2B2 Uzuner et al.
 (2011) which includes test, treatment, and problem entities, NCBI-disease Doğan et al.
@@
 (2021) and DICE Ma et al.
-(2022) formats for the biomedical NER.
+(2022) for biomedical NER.
 In TANL, the task is framed as a translation task which involves augmenting the text by tagging entity types for each word directly within the text.
@@
 Figure 1: TANL input/output format for NER task.
-Then, the generated output is decoded into the BIO format for the assessment.
+Then, the generated output is decoded into the BIO format Ramshaw and Marcus (1999) for the assessment.
 In the refined DICE format, the input-output format involves adding a description for each entity type in a template following DEGREE Hsu et al.
@@
 For the entity types with no matched entities in the sentence, the output returns <entity> token in the output.
-Examples for the NCBI-disease and BC2GM datasets are presented in Appendix A.
+Examples for the NCBI-disease and BC2GM datasets are presented in Appendix A.5.
 […]
 Figure 2: DICE input/output format for NER task.
-Our experiments in Table 1 reveal that neither format consistently outperforms the other, rather the effectiveness of each format varies depending on the complexity of the dataset and model size.
+Our experiments in Table 1 reveal that neither format consistently outperforms the other; rather, the effectiveness of each format varies depending on the complexity of the dataset and model size.
 To maintain consistency in the rest of our experiments, we opted for the TANL format, in which the input-output relationship exhibits a more straightforward pattern.
-input-output format I2B2 NCBI-disease BC2GM
-GPT-3.5-turbo DICE 50.0 62.0 55.6
-TANL 59.7 51.3 50.8
-GPT-4 DICE 70.1 77.8 67.9
-TANL 73.5 70.0 69.6
-Table 1: Comparison of TANL and DICE using GPT-3.5-turbo/GPT-4 on three benchmark datasets. The superiority of any single format varies with the complexity of the dataset and model size.
+Model input-output format I2B2 NCBI-disease BC2GM
+M/T M/T M/T
+GPT-3.5-turbo DICE 41.2 /50.0 45.3 /62.0 43.3 /55.6
+TANL 52.9/59.7 46.5/51.3 39.1/50.8
+GPT-4 DICE 58.8/70.1 68.1/77.8 57.1/67.9
+TANL 61.9/73.5 67.5/70.0 56.4/69.6
+Table 1: TANL vs. DICE format with GPT-3.5-turbo/GPT-4 . The superiority of any single format varies with the complexity of the dataset and model size.
 4 In-Context Examples Selection: A Key to Improving ICL Outcomes
-In-context examples can be randomly chosen from the training set, however, researchers have demonstrated that the performance of ICL depends on the order and similarity of ICL examples to the test samples Liu et al.
+In-context examples can be randomly chosen from the training set; however, researchers have demonstrated that the performance of ICL depends on the order and similarity of ICL examples to the test samples Liu et al.
 (2021); Min et al.
 (2022); Gao et al.
-(2023a); Zhang et al.
-(2023).
+(2023a).
 Liu et al.
@@
 We used MPNET Song et al.
-(2020) for its popularity and performance on sentence embedding benchmarks111 https://www.sbert.net/docs/pretrained_models.html, SimCSE Gao et al.
+(2020) for its popularity and performance on sentence embedding benchmarks Reimers and Gurevych (2019) , SimCSE Gao et al.
 (2021) for its documented performance as an alternative to standard sentence transformers, and BioClinicalBERT Alsentzer et al.
@@
 BioClinicalRoBERTa achieved the best results among all example encoders tested.
-The strong performance of BioClinicalBERT and BioClinicalRoBERTa underscores the importance of using LM encoders pretrained on the biomedical text when applying KATE for the biomedical NER.
-KATE method vs random sampling I2B2 NCBI-disease BC2GM
-GPT-3.5-turbo Random sampling 59.7 51.3 50.8
+The strong performance of BioClinicalBERT and BioClinicalRoBERTa underscores the importance of using LM encoders pretrained on biomedical text when applying KATE for biomedical NER.
+Model KATE vs RS I2B2 NCBI-disease BC2GM
+M/T M/T M/T
+GPT-3.5-turbo (ICL) RS 52.9/59.7 46.6/51.3 39.1/50.8
 […]
-BioClinicalBERT 78.9 78.8 72.0
+BioClinicalBERT 67.0/78.9 67.6/78.8 60.9/72.0
 […]
-SimCSE 76.1 76.1 68.8
-Hu et al. (2024b) 49.3 - -
-GPT4 Random sampling 73.5 70.0 69.6
+SimCSE 65.2/76.1 61.6/76.1 57.8/68.8
+Hu et al. (2024b) 49.3/ - - -
+GPT4 (ICL) RS 67.7/73.5 62.6/70.0 59.2/69.6
 […]
-BioClinicalBERT 88.1 88.0 79.4
+BioClinicalBERT 81.7/88.1 79.3/88.0 71.9/79.4
 […]
-SimCSE 86.6 86.5 77.9
-Hu et al. (2024b) 59.3 - -
-Table 2: Random example selection versus KATE with medical/non-medical encoders.
-KATE significantly outperforms random sampling in all settings, and LMs pretrained on the biomedical text outperform strong, general domain encoders.
+SimCSE 79.6/86.6 77.3/86.5 69.9/77.9
+Hu et al. (2024b) 59.3/ - - -
+BioBERT fully supervised - /87.3 - /89.1 - /83.8
+BioClinicBERT fully supervised - /87.7 - /89.0 - /81.7
+BioClinicRoBERTa fully supervised - /89.7 - /89.0 - /87.0
+Table 2: 16-shot ICL for Random example selection (RS) vs.
+KATE method Vs MLMs with Mention/Token-level (M/T) analysis.
+KATE significantly outperforms random sampling in all settings, and LMs pre-trained on biomedical text outperform general domain encoders.
 […]
@@
 Within the scope of LLMs for biomedical applications, an essential question is whether to prompt a closed-source LLM via ICL or fine-tune an open-source one.
-On the other hand, comparing two different LLMs employing divergent strategies is not straightforward.
-To provide some insight into this dilemma, we examined two key factors, performance and cost, for clinical NER, and present a detailed analysis under various experiment settings.
+Comparing two different LLMs employing divergent strategies is not straightforward.
+To provide some insight into this dilemma, we examined two key factors, performance and cost, for biomedical NER, and present a detailed analysis under various experiment settings.
 This comparison offers valuable perspective into the right strategy given the task and dataset attributes.
 For fine-tuning, we used LoRA Hu et al.
-(2021), and its details can be found in Appendix A.
-The cost of fine-tuning comes from training an LLM on a large labeled dataset while the cost of ICL mainly comes from calling the API for each input query.
-For 16-shot ICL experiments, we calculated the cost based on the number of processed and generated tokens considering the average text size based on the current LLM API pricing.222 https://openai.com/pricing The estimated cost for the entire test set of each benchmark dataset considering the input text, prompt, and generated text size using the TANL format is summarized in Table 3.
+(2021).
+Details can be found in Appendix A.5.
+The cost of fine-tuning comes from training an LLM on a large labeled dataset while the cost of ICL mainly comes from calling an API for each input query.
+For 16-shot ICL experiments, we calculated the cost based on the number of processed and generated tokens considering the average text size based on current LLM API pricing.111 https://openai.com/pricing The estimated cost for the entire test set of each benchmark dataset considering the input text, prompt, and generated text size using the TANL format is summarized in Table 3.
 Referring to the OpenAI API for fine-tuning pricing, we also estimated the cost for fine-tuning LLama2-7B which is summarized in Table 3.
 Interestingly, for the I2B2 dataset, GPT-3.5-turbo with a much cheaper cost outperforms fine-tuning Llama2-7B.
-Language model I2B2 NCBI-disease BC2GM
-Performance GPT-3.5-turbo w/ KATE 78.9 78.8 72.5
-GPT4 w/ KATE 88.4 88.3 80.7
-Llama2-7B 76.2 91.3 75.1
-Cost GPT3.5-turbo w/ KATE \$0.35 \$0.11 \$1.34
-GPT4 w/ KATE \$10.42 \$3.12 \$40.13
-Llama2-7B \$5.23 \$5.24 \$5.69
-Table 3: Analysis of ICL vs fine-tuning LLMs: assessing performance and cost implications.
+Model I2B2 NCBI-disease BC2GM
+M/T M/T M/T
+Performance GPT-3.5-turbo w/ KATE 67.0/78.9 68.0/78.8 61.6/72.5
+GPT4 w/ KATE 81.7/88.4 79.3/88.3 72.4/80.7
+Llama2-7B 61.2/76.2 80.4/91.3 68.1/75.1
+Cost (T+I) GPT3.5-turbo w/ KATE (\$0.35) (\$0.11) (\$1.34)
+GPT4 w/ KATE (\$10.42) (\$3.12) (\$40.13)
+Llama2-7B (\$47.85+\$7.4) (\$23.5+\$1.2) (\$69.7+\$12.9)
+Table 3: Analysis of ICL vs fine-tuning LLMs: assessing performance and cost (Training + Inference) implications.
 Fine-tuning LLama2 exhibits superior outcomes on NCBI-disease, whereas GPT-4, enhanced by KATE using a biomedical encoder, achieves more favorable results on both the I2B2 and BC2GM datasets.
@@
 This method involves retrieving relevant documents from a large corpus and providing this external knowledge in the input context to improve the quality and relevance of the generated text.
-Inspired by RAG, we developed a new method, DiRAG, to utilize UMLS as an external resource to augment the input data for the clinical NER task.
-The process with detailed prompts is visualized in Fig 3.
-Unlike traditional RAG techniques that rely on embedding similarities to retrieve relevant documents, our approach initially employs GPT-3.5-turbo to tackle a more straightforward task: identifying all words within an input text that could potentially qualify as named entities.
-Then, we look up each selected word in an external knowledge base, e.g., UMLS to augment the input data with useful information such as definition.
-Finally, we call GPT-4 with augmented input text.
-A similar prompting format to UniNER is used to do a fair comparison.
-We tested this approach on zero-shot NER and compared it with baselines in Table 4.
-Our proposed strategy with GPT-4 outperforms the baseline on I2B2 and NCBI-disease datasets with an F1 score improvement of 5.6% and 0.5% respectively.
-Additionally, the approach enhanced the performance of GPT-3.5-turbo on both the I2B2 and NCBI-disease datasets by 10.4% and 6.9% respectively.
-It proved ineffective for the BC2GM dataset due to the nature of the UMLS knowledge base which is predominantly tailored to medical terminology rather than biogenetics.
-We expect our approach to outperform GPT-4 on BC2GM if a more relevant knowledge base is employed.
-I2B2 NCBI-disease BC2GM
-UniversalNER Zhou et al. (2023) 40.4 61.4 47.2
-Rohanian et al. (2023) w/ GPT-3.5 - 33.4 32.0
-Hu et al. (2024b) w/ GPT-3.5-turbo 43.6 - -
-Hu et al. (2024b) w/ GPT-4 52.6 - -
+Inspired by RAG, we developed a new method, DiRAG, to utilize UMLS as an external resource to augment the input data for the biomedical NER task.
+The process with detailed prompts is visualized in Fig 3, while an expanded view of the UMLS component is depicted in Fig 8.
+Unlike traditional RAG techniques that rely on embedding similarities to retrieve relevant documents, our approach initially employs the LLM to tackle a more straightforward task: identifying all words that could potentially qualify as medical named entities.
+Then, we look up each selected word in an external knowledge base, e.g., UMLS to augment the input data with useful information such as term definition.
+Then, we call the LLM with augmented input text.
+The process is visualized in Fig 8.
+We tested the approach on zero-shot NER and compared it with SOTA in Table 4.
+Our proposed approach enhanced the performance of both GPT versions on the I2B2 and NCBI-disease datasets significantly.
+DiRAG with GPT-4 achieved SOTA for zero-shot NER.
+Our approach proved ineffective for the BC2GM dataset due to the nature of the UMLS knowledge base which is predominantly tailored to medical terminology rather than biogenetics.
+We expect our approach to outperform GPT-4 on BC2GM with a more relevant knowledge base.
+Model I2B2 NCBI-disease BC2GM
+M/T M/T M/T
+UniversalNER Zhou et al. (2023) 40.4/ - 60.4/ - 47.2/ -
+Rohanian et al. (2023) w/ GPT-3.5 - 33.4 / - 32.0 / -
+Hu et al. (2024b) w/ GPT-3.5-turbo 39.3/ - - -
+Hu et al. (2024b) w/ GPT-4 52.6/ - - -
 […]
-GPT-3.5-turbo w/ DiRAG 43.55 45.0 28.8
+GPT-3.5-turbo w/ DiRAG 43.0 / 55.7 44.7 / 50.0 30.45 / 22.5
 […]
-GPT-4 w/ DiRAG 60.04 62.17 53.83
-Table 4: Zero-shot clinical NER using GPT-3.5-turbo and GPT-4 w/ and w/o DiRAG in comparison to SOTA. Our method improved zero-shot NER significantly for I2B2 and NCBI-disease datasets.
+GPT-4 w/ DiRAG 53.1 /62.8 61.0 /66.2 51.1 / 55.0
+Table 4: Zero-shot NER with GPT models w/ and w/o DiRAG vs.
+SOTA.
+DiRAG improved zero-shot NER significantly for I2B2 and NCBI-disease datasets for both GPT models.
+Results with confidence intervals are in the appendix.
 7 Conclusion
-In this paper, we explored LLMs by customizing various prompting techniques specifically for the biomedical NER.
-Through a detailed comparative analysis, we have highlighted the vital role of ICL and the selection of contextually pertinent examples, demonstrating the benefits to clinical NER tasks.
-Moreover, our investigation into incorporating external clinical knowledge resulted in a novel data augmentation approach, considerably advancing the capabilities of zero-shot clinical NER with LLMs.
+We explored LLMs for biomedical NER by customizing various prompting techniques.
+Through a detailed comparative analysis, we highlighted the vital role of ICL and the selection of contextually pertinent examples with biomedical text encoders for biomedical NER tasks.
+Moreover, our investigation into incorporating external medical knowledge resulted in a novel data augmentation approach, considerably advancing the capabilities of zero-shot biomedical NER with LLMs.
 Limitations
-While we have shown the potential of enhancing LLM performance for clinical NER, the experiments in this paper are limited in two aspects mainly due to computational constraints.
+While we have shown the potential of enhancing LLM performance for biomedical NER, the experiments in this paper are limited in two aspects mainly due to computational constraints.
 (1) TANL uses a straightforward text-to-text format while DICE uses additional descriptions.
@@
 Ablation studies on components of each format could help researchers design new prompt formatting strategies.
-(2) Our RAG-based method exclusively utilizes UMLS as the knowledge base which is limited in its vocabulary.
-For medical terms not covered by UMLS, we did not augment the input text, though other knowledge bases such as Wikipedia could potentially serve as an alternative.
-8 Aknowledgment
-This work was supported in part by Optum. We would like to thank the reviewers for their feedback to improve this research work.
+(2) Our RAG-based method exclusively utilizes UMLS as the knowledge base, though it is limited in its vocabulary.
+For medical terms not covered by UMLS, we did not augment the input text.
+Other knowledge bases such as Wikipedia could serve as an alternative.
 Appendix A Appendix
-Fine-tuning setting for Llama
-We fine-tuned Llama2-7B for three epochs and maintained a batch size of 16, learning rate of 2e-4, and cap the maximum sequence length at 512, truncating any sequences that exceeded this limit.
+A.1 TANL/DICE more examples
+In Fig 4-8, we visualize two examples of each format for NCBI-disease and BC2GM datasets for more demonstration.
+A.2 Benchmark datasets
+We studied LLMs on I2b2, NCBI-disease, and BC2GM dataset. In the following, we provide some details about each.
+I2B2: I2B2 is a collection of annotated clinical records that are used primarily for Clinical NER.The task involves identifying clinical terms such as medical problems, treatments, and tests from patient records.
+The dataset typically includes a large number of annotated clinical narratives that are de-identified to protect patient confidentiality.
+This makes it a rich resource for training and testing NER models.
+NCBI-disease: This dataset is specifically curated for disease name recognition and normalization in biomedical texts.
+It comprises abstracts from PubMed annotated for disease mentions and linked to the NCBI disease database.
+The corpus is relatively smaller compared to i2b2 but is densely annotated, providing high-quality, fine-grained annotations of disease entities, which are crucial for models aimed at medical literature.
+BC2GM: This dataset focuses on the recognition of gene and gene product mentions in PubMed abstracts that is a suitable dataset for biological NER.
+The BC2GM dataset is extensively annotated to include a wide range of gene and gene product mentions, reflecting the complex and varied ways these entities are referred to in scientific literature.
+A.3 PEFT setting of Llama for fine-tuning
+We fine-tuned Llama2-7B on the entire training set of each dataset for three epochs and maintained a batch size of 16, learning rate of 2e-4, and cap the maximum sequence length at 512, truncating any sequences that exceeded this limit.
 The LoRA dropout rate is adjusted to 0.1, and the LoRA $\alpha$ and rank parameters are also set at 16 and 32 respectively.
-The training was done on 4 NVIDIA Tesla V100 GPUs for an approximate of 2 days.
-TANL and DICE more examples
-In this section, we visualize two examples of each format for NCBI-disease and BC2GM datasets for demonstration.
+The training was done on 4 NVIDIA Tesla V100 GPUs for approximately 24, 12, and 63 hours for I2B2, NCBI-disease, and BC2GM respectively.
+A.4 Few-shot and Zero-shot performances with Confidence Interval
+We introduced both few-shot and zero-shot settings to comprehensively evaluate the versatility and generalization capabilities of our study across different levels of data availability.
+While it’s true that the performance in the zero-shot setting is generally lower compared to the few-shot setting, this approach offers valuable insights into the model’s behavior when no training examples are provided.
+The zero-shot setting, leveraging techniques like Retrieval-Augmented Generation (RAG), demonstrates the model’s potential to utilize pre-existing knowledge embedded in its parameters and external sources effectively.
+This is particularly important for scenarios where labeled data is scarce or unavailable, making zero-shot learning a critical area of study to ensure broader applicability of the model in real-world applications.
+Moreover, the inclusion of both methodologies allows us to highlight the performance trade-offs and strengths of the model under different instructional paradigms, contributing to a more robust and nuanced understanding of its capabilities.
+We ran all experiments with different random seeds and reported the full results of Table 2, 3, and 4 with confidence Intervals In Tables 6, 7, and 8.
+A.5 UMLS detail
+In Fig 8, we visualize the process by which potential words suggested by the LLM are searched within the UMLS and demonstrate how the input is augmented to enhance zero-shot prompting in LLMs.
 […]
@@
 Figure 7: DICE input-output format example for BC2GM dataset
+Mention/Token
+Model input-output format I2B2 NCBI-disease BC2GM
+GPT-3.5-turbo DICE 41.2 $\pm$ 0.2 /50.0 $\pm$ 0.1 45.3 $\pm$ 0.2 /62.0 $\pm$ 0.3 43.3 $\pm$ 0.5 /55.6 $\pm$ 0.4
+TANL 52.9 $\pm$ 0.3/59.7 $\pm$ 0.4 46.5 $\pm$ 0.5/51.3 $\pm$ 0.5 39.1 $\pm$ 0.4/50.8 $\pm$ 0.5
+GPT-4 DICE 58.8 $\pm$ 0.4 /70.1 $\pm$ 0.3 68.1 $\pm$ 0.9/77.8 $\pm$ 1.1 57.1 $\pm$ 0.6/67.9 $\pm$ 0.5
+TANL 61.9 $\pm$ 0.3/73.5 $\pm$ 0.5 67.5 $\pm$ 0.8/70.0 $\pm$ 0.6 56.4 $\pm$ 0.2/69.6 $\pm$ 0.3
+Table 5: TANL vs. DICE format with GPT-3.5-turbo/GPT-4 with confidence intervals
+Model KATE vs RS I2B2 NCBI-disease BC2GM
+M/T M/T M/T
+GPT-3.5-turbo (ICL) RS 52.9 $\pm$ 0.3 / 59.7 $\pm$ 0.4 46.6 $\pm$ 0.5 / 51.3 $\pm$ 0.5 39.1 $\pm$ 0.4 / 50.8 $\pm$ 0.5
+[…]
+BioClinicalBERT 67.0 $\pm$ 0.6/78.9 $\pm$ 0.5 67.6 $\pm$ 0.1/78.8$\pm$ 0.1 60.9 $\pm$ 0.7 /72.0 $\pm$ 0.5
+[…]
+SimCSE 65.2 $\pm$ 0.2 / 76.1 $\pm$ 0.3 61.6 $\pm$ 0.4 / 76.1 $\pm$ 0.3 57.8 $\pm$ 0.5 / 68.8 $\pm$ 0.4
+Hu et al. (2024b) 49.3/ - - -
+GPT4 (ICL) RS 67.7 $\pm$ 0.3 / 73.5 $\pm$ 0.5 62.6 $\pm$ 0.8 /70.0 $\pm$ 0.6 59.2 $\pm$ 0.2 / 69.6 $\pm$ 0.3
+[…]
+BioClinicalBERT 81.7 $\pm$ 0.4/88.1 $\pm$ 0.3 79.3 $\pm$ 0.4/88.0 $\pm$ 0.3 71.9 $\pm$ 0.3 / 79.4 $\pm$ 0.3
+[…]
+SimCSE 79.6 $\pm$ 0.5 / 86.6 $\pm$ 0.4 77.3 $\pm$ 0.5 / 86.5 $\pm$ 0.8 69.9 $\pm$ 0.8 / 77.9 $\pm$ 0.5
+Hu et al. (2024b) 59.3/ - - -
+BioBERT fully supervised - /87.3 - /89.1 - /83.8
+BioClinicBERT fully supervised - /87.7 - /89.0 - /81.7
+BioClinicRoBERTa fully supervised - /89.7 - /89.0 - /87.0
+Table 6: Random example selection (RS) vs.
+KATE with medical/non-medical encoders vs.
+fully supervised models with Mention/Token-level (M/T) analysis.
+KATE significantly outperforms random sampling in all settings, and LMs pre-trained on the biomedical text outperform strong, general domain encoders.
+HunFlair is added to the paper
+Model I2B2 NCBI-disease BC2GM
+M/T M/T M/T
+Performance GPT-3.5-turbo w/ KATE 67.0 $\pm$ 0.6 / 78.9 $\pm$ 0.5 68.0 $\pm$ 0.3 / 78.8 $\pm$ 0.1 61.6 $\pm$ 0.1 / 72.5 $\pm$ 0.6
+GPT4 w/ KATE 81.7 $\pm$ 0.4 / 88.4 $\pm$ 0.6 79.3 $\pm$ 0.4 / 88.3 $\pm$ 0.8 72.4 $\pm$ 0.6 / 80.7 $\pm$ 0.4
+Llama2-7B 61.2 $\pm$ 1.8 / 76.2 $\pm$ 1.3 80.4 $\pm$ 0.9/ 91.3 $\pm$ 1.1 68.1 $\pm$ 1.4 / 75.1 $\pm$ 1.3
+Cost (T+I) GPT3.5-turbo w/ KATE (\$0.35) (\$0.11) (\$1.34)
+GPT4 w/ KATE (\$10.42) (\$3.12) (\$40.13)
+Llama2-7B (\$47.85+\$7.4) (\$23.5+\$1.2) (\$69.7+\$12.9)
+Table 7: Analysis of ICL vs fine-tuning LLMs: assessing performance and cost implications.
+Fine-tuning LLama2 exhibits superior outcomes on NCBI-disease, whereas GPT-4, enhanced by KATE using a biomedical encoder, achieves more favorable results on both the I2B2 and BC2GM datasets.
+Model I2B2 NCBI-disease BC2GM
+M/T M/T M/T
+UniversalNER Zhou et al. (2023) 40.4/ - 60.4/ - 47.2/ -
+Rohanian et al. (2023) w/ GPT-3.5 - 33.4 / - 32.0 / -
+Hu et al. (2024b) w/ GPT-3.5-turbo 39.3/ - - -
+Hu et al. (2024b) w/ GPT-4 52.6/ - - -
+HunFlair Weber et al. (2021) 0.0 / 0.0 24.8 / 36.1 28.2 / 22.7
+[…]
+GPT-3.5-turbo w/ DiRAG 43.0 $\pm$ 0.9 / 55.7 $\pm$ 1.5 44.7 $\pm$ 0.5 / 50.0 $\pm$ 2.1 30.45 $\pm$ 1.6 / 22.5 $\pm$ 2.1
+[…]
+GPT-4 w/ DiRAG 53.1 $\pm$ 1.1 / 62.8 $\pm$ 1.2 61.0 $\pm$ 0.6 /66.2 $\pm$ 0.5 51.1 $\pm$ 2.0 / 55.0 $\pm$ 2.2
+Table 8: Full results of Zero-shot NER with GPT-3.5-turbo and GPT-4, w/ and w/o DiRAG, and their comparision with SOTA.
+Our method improved zero-shot NER significantly for I2B2 and NCBI-disease datasets.
+[…]
+Figure 8: UMLS search.
+The GPT model is prompted for a simpler task of identifying all words that could potentially be a named entity.
+Then, the retrieved information from UMLS will augment the original input text for recalling the LLM